OpenStudy (anonymous):

What is the slope of the line that contains the points (-2, 5) and (7, 5)? A. -9 B. 0 C. 9/10 D. -9/10

OpenStudy (anonymous):

Would I add the answers like I did with the last one?

OpenStudy (anonymous):

M= y1 - y2/ x1-x2

OpenStudy (arnavguddu):

for given points (x1, y1) and (x2,y2)..thru which a line pass slope = (y2-y1)/(x2-x1) = (difference of y values) / (diffof x values) ... in the same order

OpenStudy (anonymous):

ok correct me if i did this wrong?